OpenStudy (marissalovescats):

@jim_thompson5910 I have some problems here I wouldn't mind you checking/the 3rd one helping me answer them. That'd be swell :)

OpenStudy (marissalovescats):

Just let me attach a file in a second

jimthompson5910 (jim_thompson5910):

#1 is correct

OpenStudy (marissalovescats):

For number 3 the answer choices are: f(0)=2 f(x) does not equal 2 for all x>=0 f(2) is undefined lim x approaches 2 f(x)= infinity lim x approaches infinity f(x)=2 I think It's the 3rd one but I'm not fully convinced in myself.

jimthompson5910 (jim_thompson5910):

#2 is a bit off the integral of e^(-4x) is NOT -4e^(-4x)

OpenStudy (marissalovescats):

Oh darn then what is it?

jimthompson5910 (jim_thompson5910):

it's (-1/4)*e^(-4x)

jimthompson5910 (jim_thompson5910):

you can check by deriving (-1/4)*e^(-4x) y = (-1/4)*e^(-4x) y' = d/dx[ (-1/4)*e^(-4x)] y' = -4*(-1/4)e^(-4x) y' = e^(-4x)

OpenStudy (marissalovescats):

Oh right is it because you set u=-4x and du=-4dx so -1/4du=dx?

jimthompson5910 (jim_thompson5910):

To integrate e^(-4x), I'd let u = -4x so du = -4dx ---> dx = -du/4 ---> dx = (-1/4)dx

jimthompson5910 (jim_thompson5910):

exactly

OpenStudy (marissalovescats):

So it's 1/4-e^-4/4

jimthompson5910 (jim_thompson5910):

one sec

jimthompson5910 (jim_thompson5910):

If you meant \[\Large \frac{1}{4} - \frac{1}{4}e^{-4}\] then you are correct

OpenStudy (marissalovescats):

Well it's the same thing as \[\frac{ 1 }{ 4 }-\frac{ e ^{-4} }{ 4} \]

jimthompson5910 (jim_thompson5910):

correct

OpenStudy (marissalovescats):

Cool and for the 3rd one? I think it's the 3rd answer but not fully sure

jimthompson5910 (jim_thompson5910):

I see the answer choices, but does it say what the function f is?

jimthompson5910 (jim_thompson5910):

oh nvm, we don't need the function lol

jimthompson5910 (jim_thompson5910):

f(2) may or may not be defined, it's hard to say but all we know is that the horizontal asymptote is y = 2

jimthompson5910 (jim_thompson5910):

so if you the horizontal asymptote is y = 2, what can you say for sure about f(x)?

OpenStudy (marissalovescats):

Uhmmm I think that f(x) cannot equal 2 for all x>=0?

OpenStudy (marissalovescats):

WAIT no no no

jimthompson5910 (jim_thompson5910):

a function/graph can cross the horizontal asymptote

OpenStudy (marissalovescats):

It's an asymptote so the limit as x approaches 2 f(x)= infinity

jimthompson5910 (jim_thompson5910):

you just described the vertical asymptote at x = 2

OpenStudy (marissalovescats):

Oh darn I guess I wasn't visualizing (I was bad at the limits chapter unfortunately) So is it f(0)=2 because it's a horizontal line and they have a slope of 0? I'm not sure to be honest.

jimthompson5910 (jim_thompson5910):

you were on the right track when you brought up the limit

OpenStudy (marissalovescats):

Oh I couldn't visualize the other limit so I just guessed :P

jimthompson5910 (jim_thompson5910):

as x approaches infinity, then f(x) will approach 2 it could cross over the asymptote (as many times as you want), but overall, as x gets really huge, f(x) will settle on 2

jimthompson5910 (jim_thompson5910):

and we write that as \[\Large \lim_{x\to \infty} f(x) = 2\]

OpenStudy (marissalovescats):

Oh right is that the thing like this 

jimthompson5910 (jim_thompson5910):

yeah that's one of many ways to visualize it

OpenStudy (marissalovescats):

Oh alright I understand now

jimthompson5910 (jim_thompson5910):

that's good

OpenStudy (marissalovescats):

Can you check the first two and help me with the 3rd? I suck at limits And in the second one it's rad2/2-1 I just forgot to rewrite the 2 in the denominator because it looked messy

jimthompson5910 (jim_thompson5910):

#4 is correct. Nice job

jimthompson5910 (jim_thompson5910):

for #5, the integral of sin(x) is -cos(x) not just cos(x)

jimthompson5910 (jim_thompson5910):

this is because if y = cos(x), then dy/dx = -sin(x) so we get close to what we want, we're just off by a sign if we have y = -cos(x), then dy/dx = -(-sin(x)) = sin(x)

OpenStudy (marissalovescats):

Oh right

OpenStudy (marissalovescats):

So -rad2/2+1

jimthompson5910 (jim_thompson5910):

good

OpenStudy (marissalovescats):

And for the 3rd I think I have to do long division then solve for x but i dont remember how to do long division for polynomials lol

jimthompson5910 (jim_thompson5910):

you could do long division, but it would take a very long time and there's a potential for error

jimthompson5910 (jim_thompson5910):

for #6, you can divide everything by x^3 to go from \[\Large \frac{x^3 - 2x^2+3x-4}{4x^3 - 3x^2+2x-1}\] to \[\Large \frac{\frac{x^3}{x^3} - \frac{2x^2}{x^3}+\frac{3x}{x^3}-\frac{4}{x^3}}{\frac{4x^3}{x^3} - \frac{3x^2}{x^3}+\frac{2x}{x^3}-\frac{1}{x^3}}\] and that turns into \[\Large \frac{1 - \frac{2}{x}+\frac{3}{x^2}-\frac{4}{x^3}}{4 - \frac{3}{x}+\frac{2}{x^2}-\frac{1}{x^3}}\]

OpenStudy (marissalovescats):

Why would I divide everything by x^3? That looks really unfamiliar to me he always has us to long division and I never remember lol

jimthompson5910 (jim_thompson5910):

like I said, you could do long division, but be cause the denominator is so large, it'd take a while

jimthompson5910 (jim_thompson5910):

I'm dividing by x^3 since it's the largest term with the largest exponent

jimthompson5910 (jim_thompson5910):

doing that will turn each term into a mini fraction (within the fraction)

jimthompson5910 (jim_thompson5910):

each mini fraction will then go to 0 as x --> infinity

OpenStudy (marissalovescats):

Okay and then what would you do after that?

jimthompson5910 (jim_thompson5910):

so it effectively eliminates a lot of the terms

jimthompson5910 (jim_thompson5910):

\[\Large \frac{1 - \frac{2}{x}+\frac{3}{x^2}-\frac{4}{x^3}}{4 - \frac{3}{x}+\frac{2}{x^2}-\frac{1}{x^3}}\] turns into \[\Large \frac{1 - 0 + 0 - 0}{4 - 0 + 0 - 0}\] when x approaches infinity

OpenStudy (marissalovescats):

What how?

jimthompson5910 (jim_thompson5910):

for instance \[\Large \lim_{x \to \infty} \left(\frac{2}{x}\right) = 0\]

jimthompson5910 (jim_thompson5910):

each mini fraction will have the same limit...0

OpenStudy (marissalovescats):

Oh right then I don't even have to do the divide by x^3, I can just plug in 0 and get 4 can't I? So wait idk lol

jimthompson5910 (jim_thompson5910):

not quite

jimthompson5910 (jim_thompson5910):

remember x is approaching infinity, not 0